Q1) Find the Eigenvalues and the Eigenvectors for the following matrices16
Answer #1

-i1 A= -10 |-11-> -lo-0 (9-入エ) - (-0 (-1-X - Lo7 5-7 l1-N(5--Gx-10) - -55+11-5t+6 - +65 =E =X0tT(At57 - Ots) (At - 1- 0 (A-AIR2 1ORTER -10 - 10 1091 + (-10)0 Lef T2-k - K tK Vechr FI-(-) (A-AI)= ( 5-(-5) -6 |a ReE RtR -63-10 =*1 97 S cifaic is V k Яplease rate my answer